# TalkSketch: Multimodal Generative AI for Real-time Sketch Ideation with Speech

### v1

@@ Abstract
Introduction
During early-stage design, sketching plays a central role as an improvisational, open-ended, and dynamic practice (landay1995interactive; suwa2022roles). Designers frequently switch between design phases and iterate their sketches to explore alternatives and refine ideas toward promising directions (buxton2010sketching). To enhance efficiency in this formative stage, researchers have explored integrating generative AI (GenAI) (fui2023generative) into sketch-based design workflows (zhang2023generative; davis2025sketchai; lin2025inkspire; zhou2025partstickers). With advances in multimodal large language models (LLMs), GenAI now offers stronger creative capabilities (vinker2025sketchagent; ccelen2024design; xu2024llm), making it increasingly suitable for supporting designers in early-stage ideation.
However, despite the growing use of GenAI chatbots in idea generation, designers often struggle to craft effective prompts and to express evolving visual concepts through text alone. This command-based interaction paradigm (subramonyam2024bridging) places the burden on designers to direct the system, overlooking other forms of input that may be more natural and intuitive during creative work (kim2009study; purcell1998drawings). In practice, designers often verbalise ideas spontaneously while sketching, yet such contextual cues are rarely captured by current systems.
@@ Introduction
Related Work
Multimodal Human–AI Interaction for Creative Ideation
Conversational User Interfaces (CUIs) enable dialogue-based interaction that mimics human conversation (mctear2002spoken) and are widely deployed in chatbots (folstad2017chatbots) and voice-activated assistants such as Amazon Alexa and Apple’s Siri (de2020intelligent). Recent advances in Generative AI (GenAI) (feu2024generative; tank2024metacognitive), including large language models (LLMs) such as ChatGPT (achiam2023gpt) and multimodal assistants such as Gemini (team2024gemini), have made conversational human–AI interaction broadly accessible and flexible. These tools are increasingly applied in domains such as education (shaer2024ai), healthcare (hedderich2024piece), and creative work (angert2023spellburst; zhang2025breaking), supporting users through natural language dialogue.
A central concern in these systems is whether users can express intent clearly and efficiently. Prior work has explored multimodal input and interface strategies to improve intent communication. Hu et al. (hu2025gesprompt) introduced GesPrompt, which combines co-speech gestures with voice input to enable more natural intent expression in Extended Reality (XR) environments. Cho et al. (cho2025persistent) proposed Persistent Assistant, integrating embodied input and multimodal feedback for seamless everyday interactions. Other work embeds LLMs into direct manipulation environments, mapping GUI actions or visual edits into structured prompts (masson2024directgpt; angert2023spellburst), showing that embedding language models within interactive contexts can reduce the burden of verbose prompting. Yet, most of these systems are designed for short, directive, or domain-specific tasks such as code editing or data visualisation, rather than open-ended creative ideation.
In design, creativity-support tools increasingly integrate GenAI into mainstream workflows. Commercial platforms such as Adobe Firefly (adobefirefly2025tools), Canva Magic Studio (canva2025overview), and Figma AI (figmaai2025update) demonstrate how generative models assist in rapid visual exploration, while research prototypes extend this trend toward more expressive input. For example, DesignPrompt (peng2024designprompt) allows designers to compose prompts using text, colour, and imagery, and DesignWeaver (tao2025designweaver) introduces palette-based refinement to support iterative design. Similarly, sketch-based interfaces such as Inkspire (lin2025inkspire) and SketchAI (davis2025sketchai) demonstrate how freehand input can guide image generation and support analogical inspiration.
@@
Talking as a Novel Input Modality for Sketching
Speech is also a natural modality for externalising thought, long recognised in systems such as SHRDLU (winograd1972understanding) and Put-That-There (bolt1980put) from the 1970-80s. It enables real-time, low-friction expression of ideas, making it especially valuable in early-stage design when thoughts are fluid and evolving (fry1979physics). Sketching and speech together can serve as two intuitive, complementary interaction methods in design and creative domains, allowing users to express nuances that one modality alone might omit  (adler2007speech).
Recent works demonstrate how combining sketch and speech supports more natural and expressive communication: Rosenberg et al. (rosenberg2024drawtalking) introduced DrawTalking, which combines sketching with storytelling through speech to construct interactive animated worlds, while Giunchi et al. (giunchi2021mixing) showed that integrating sketch and speech for 3D model retrieval in virtual environments helped overcome the limitations of sketch-only input. Cheng et al. (cheng2025aiawareness) also explored speech and sketch as inputs for enhancing communication with AI through context awareness in an exploratory Wizard-of-Oz study.
@@
Formative Study: Understanding Early-Stage Design Workflow with GenAI
We first conducted a formative study, combining a design task with interviews, to understand how designers integrate GenAI into early-stage design and to identify opportunities for multimodal AI support to inform our subsequent system design in Section TalkSketch: System Design and Development. Our institution’s ethics review board approved this study [number redacted for review], and we obtained informed consent from all participants.
Participants We recruited six participants (2 female, 4 male) with a basic background in design; at the very least, they had completed one design-related course (e.g., Urban Sketching or Design Thinking and Innovation). The participants included three students, two entry-level designers, and one experienced design practitioner, with design backgrounds spanning architecture, furniture, interior, robotics, and electronic product design. All participants had 2-5 years of part/full-time experience in design. They received approximately USD 7.8 as compensation for completing the 1-hour study.
Study Protocol The study consists of three parts: pre-task interview (10 minutes), design task (30 minutes), and post-task interview (20 minutes). The pre-task interview captures participants’ design backgrounds, sketching habits, and prior experiences with GenAI tools. For the design task, we asked participants to design a household bread toaster and to express as many design ideas as possible under 30 minutes using a sketching application such as Goodnotes [^1] https://www.goodnotes.com/ or Procreate [^2] https://procreate.com/. They were allowed to use their own devices (to provide for a familiar setup) or our iPad Pro 13” and Apple Pencil set up with popular sketching apps (e.g., Goodnotes, Procreate) and GenAI tools (e.g., ChatGPT (GPT-4o) [^3] https://chatgpt.com/, Gemini [^4] https://gemini.google.com/app, Midjourney [^5] https://midjourney.com). The post-task interview then examined participants’ views on the AI’s role in supporting or limiting ideation, and captured their aspirations and expectations for an ”ideal” AI assistant.
Data Analysis We analysed the qualitative data by examining both participants’ interactions with sketching and AI tools during the design tasks and their reflections in post-study interviews. This allowed us to capture overall usage patterns, such as how participants sketched, engaged with GenAI tools, and iterated on ideas, while also identifying the challenges and gaps they experienced with current tools. These insights will then inform the design goals of our system.
Overall Usage Patterns and Challenges
@@
(1) Using GenAI for research and ideation. Participants often used ChatGPT to explore design problems, define target users, or to inspire new directions. For instance, P1 asked about commonly found toaster issues and focused on cleaning functions, while P3 explored different user groups to inspire persona-driven designs. These uses supported functional exploration, but initial AI responses were often “too generic” (P3) and only became actionable after repeated prompt refinement.
(2) Using GenAI to render sketch-based ideas. Several participants uploaded their sketches to ChatGPT Image or Gemini, or relied on text prompts, to visualise concepts. For instance, P2 uploaded a sideways-eject sketch, while P5 asked Gemini for a toaster with butter and egg compartments. However, the outputs often failed to match intent. P1 described the results as “kind of crazy,” P2 concluded, “I might have to redraw the whole thing so AI can understand,” and P5 added, “It takes too much time, I’d rather just draw.” Many participants ultimately returned to manual sketching for better clarity and control.
(3) Iterative loops across sketching, prompting, and referencing. Participants often moved back and forth between tools. P3 only used text in ChatGPT to imagine playful forms (e.g., cat- or book-shaped toasters), while P1 first researched in ChatGPT, then sketched, went back to sketching after clarifications, and finally tried rendering images. P1 noted that “uploading sketches to ChatGPT was not so easy to operate on iPad.” This reflected how tool-switching slowed the process and introduced friction, especially for those who wanted to use image generation (P1, P2, P5, P6).
Challenges across patterns. Across these usage patterns, participants encountered three main challenges: (1) AI responses that were too generic or required extensive refinement, (2) mismatched or low-quality image outputs that failed to convey intent, and (3) fragmented workflows due to frequent switching between different tools. These issues often interrupted the ideation flow and made participants rely more heavily on manual sketching for clarity and control.
Design Goals for TalkSketch
@@
Goal 1: Integrate AI with sketching tools.
@@
Goal 2: Reduce designers’ fatigue with long prompts.
@@
Goal 3: A more proactive and context-aware AI.
@@
TalkSketch: System Design and Development
@@
We first describe the technical implementation of TalkSketch (Figure 1), which integrates three core components: a digital sketching canvas for creating and managing drawings, to address the goal of integrating GenAI into sketching tools (Goal 1); a speech capture module that transcribes the user’s speech during sketching, aiming to support the goal of reducing fatigue from long prompts (Goal 2); and a multimodal AI chatbot that combines automatic insights with interactive text and image generation, aiming to make AI more proactive and context-aware (Goal 3). We then illustrate how these components work together through a system walkthrough using a concept toaster design example (Figure 2).
@@
Technical Implementations TalkSketch integrates a unified multimodal AI chatbot that aims to support both proactive and user-initiated interaction. It comprises three complementary modules: Sketching, Talking, and Multimodal AI Chatbot.
@@
Sketching Module (Figure 1 (a)) The sketching canvas (Figure 1a) serves as the central workspace where participants draw early-stage design ideas using a stylus or touch input. Built using Fabric.js [^6] http://fabricjs.com/, the canvas supports drawing, erasing, selection, undo/redo, and canvas reset as illustrated in Figure 1. The top toolbar consolidates essential drawing controls along with the Generate with AI button. To facilitate interaction between users and AI, users can select any region of the sketching canvas and export it into the chatbot as part of a multimodal prompt. The Save to Gallery button allows users to save their canvas to the gallery once they are content with it; saved canvases can also be retrieved through the Gallery.
@@
Talking Module (Figure 1 (b)) TalkSketch augments the sketching experience with real-time voice capture to support think-aloud workflows. Audio recording initiates automatically when users are sketching (without the chatbot open), and the recording symbol will be turned on simultaneously to indicate this. Once the user opens the AI Chatbot, the recording stops and the captured audio is streamed for low-latency transcription via Google Cloud Speech-to-Text [^7] https://cloud.google.com/speech-to-text.
@@
Multimodal AI Chatbot Module (Figure 1 (c) and 1 (d)) The multimodal AI chatbot (see Figures 1c and 1d) has two main components: (1) AI Insights, which automatically generate reflective feedback based on user sketches and verbal ideation, and (2) a Multimodal Chatbot Interface, where users can engage the AI through both text and sketch input. Both components share a common back-end powered by the Gemini series of multimodal models [^8] https://ai.google.dev/gemini-api/docs/models , known for their strong cross-modal reasoning capabilities and broad adoption across creative and analytical tasks (yue2024mmmu; team2024gemini). Specifically, the text-based conversational features are supported by Gemini 2.5 Flash, while image generation is handled by Gemini 2.5 Flash Image. The Multimodal AI Chatbot Module maintains a unified conversation history that persists throughout the entire session and is consistently shared across both models, ensuring coherent context when switching between text and image generation.
@@
On the back-end, the AI Insights module is powered by Gemini 2.5 Flash and customised to simulate the role of a design thinking expert. Following the Double Diamond framework (tschimmel2012design), it guides users through the Discover and Define stages by identifying potential user needs, pain points, and framing design questions, before suggesting several exploratory directions to pursue. Two complementary prompt templates were developed to support different phases of the ideation process. When a user begins drawing on a new canvas, the system automatically triggers the Kickoff Prompt to initiate ideation; in subsequent interactions, as the user continues sketching or revising their transcript, the Refine Prompt is triggered to provide iterative feedback aligned with the evolving design concept. These prompts were iteratively tested to balance interpretability, creativity support, and brevity of response.
@@
TalkSketch aims to support designers in early-stage ideation by combining freehand sketching, verbal thinking, and multimodal AI assistance. We illustrate its use through a walkthrough exemplar (Figure 2) where a fictional example user Sky is tasked with designing a household toaster. Her interaction follows a natural flow: sketching while speaking aloud, triggering the chatbot interface, reviewing automatically generated AI insights, and exploring further via multimodal interaction.
Sketching while Talking. Sky begins the design session by drawing a rough concept of a cube-shaped toaster with symbolic radial dials and a handle (see Figure 1a). As she sketches, she verbalises her thoughts, “I’m thinking of something bold and square, with a dial for heat control.” Her speech is recorded and transcribed in real time (see Figure 1b), forming a synchronised verbal-visual trace of her ideation process.
Entering the Chatbot and Viewing AI Insights. When Sky clicks Generate with AI, the system enters the chatbot interface (see Figure 1d) and immediately displays an AI Insight (see Figure 1c). This insight is automatically generated from her sketch and think-aloud transcript, offering concise design reflections based on the Double Diamond design thinking framework. The system suggests emphasising geometric form or exploring tactile dials as directions worth pursuing. Sky reads the insight and finds it helpful in framing her design direction. She realises she had forgotten to mention a few details aloud–such as her intention for the toaster to have a retractable cord–and decides to edit the transcript directly. After adding the missing information, she regenerates the AI Insight to see how the suggestions change in response to the fuller context.
Exploring via Multimodal Interaction. Sky then decides to follow up on the AI’s suggestions. She first types into the chatbot input box: “Could you give me some ideas for drawing a novel toaster?” In Text Generation Mode, the chatbot responds with directions such as transparent exteriors and interactive touchscreens. Sky incorporates several of these into her sketch. To visualise the design, she switches to Image Generation Mode, prompting: “Could you generate a realistic product based on my sketch?” The system returns a refined visual that preserves key features from her drawing.
@@
Potential Results
Enhancing Intent Expression and Interaction Naturalness

### v2

@@ Abstract
1 Introduction
During early-stage design, sketching plays a central role as an improvisational, open-ended, and dynamic practice [landay1995interactive, suwa2022roles]. Designers frequently switch between design phases and iterate their sketches to explore alternatives and refine ideas toward promising directions [buxton2010sketching]. To enhance efficiency in this formative stage, researchers have explored integrating generative AI (GenAI) [fui2023generative] into sketch-based design workflows [davis2025sketchai, lin2025inkspire, zhang2023generative, zhou2025partstickers]. With advances in multimodal large language models (LLMs), GenAI now offers stronger creative capabilities [vinker2025sketchagent, xu2024llm, ccelen2024design], making it increasingly suitable for supporting designers in early-stage ideation.
However, despite the growing use of GenAI chatbots in idea generation, designers often struggle to craft effective prompts and to express evolving visual concepts through text alone. This command-based interaction paradigm [subramonyam2024bridging] places the burden on designers to direct the system, overlooking other forms of input that may be more natural and intuitive during creative work [kim2009study, purcell1998drawings]. In practice, designers often verbalise ideas spontaneously while sketching, yet such contextual cues are rarely captured by current systems.
@@ 1 Introduction
2 Related Work
2.1 Multimodal Human–AI Interaction for Creative Ideation
Conversational User Interfaces (CUIs) enable dialogue-based interaction that mimics human conversation [mctear2002spoken] and are widely deployed in chatbots [folstad2017chatbots] and voice-activated assistants such as Amazon Alexa and Apple’s Siri [de2020intelligent]. Recent advances in Generative AI (GenAI) [feu2024generative, tank2024metacognitive], including large language models (LLMs) such as ChatGPT [achiam2023gpt] and multimodal assistants such as Gemini [team2024gemini], have made conversational human–AI interaction broadly accessible and flexible. These tools are increasingly applied in domains such as education [shaer2024ai], healthcare [hedderich2024piece], and creative work [angert2023spellburst, zhang2025breaking], supporting users through natural language dialogue.
A central concern in these systems is whether users can express intent clearly and efficiently. Prior work has explored multimodal input and interface strategies to improve intent communication. Hu et al. [hu2025gesprompt] introduced GesPrompt, which combines co-speech gestures with voice input to enable more natural intent expression in extended reality environments. Cho et al. [cho2025persistent] proposed Persistent Assistant, integrating embodied input and multimodal feedback for seamless everyday interactions. Other work embeds LLMs into direct manipulation environments, mapping graphical user interface actions or visual edits into structured prompts [angert2023spellburst, masson2024directgpt], showing that embedding language models within interactive contexts can reduce the burden of verbose prompting. Yet, most of these systems are designed for short, directive, or domain-specific tasks such as code editing or data visualisation, rather than open-ended creative ideation.
In design, creativity-support tools increasingly integrate GenAI into mainstream workflows. Commercial platforms such as Adobe Firefly [adobefirefly2025tools], Canva Magic Studio [canva2025overview], and Figma AI [figmaai2025update] demonstrate how generative models assist in rapid visual exploration, while research prototypes extend this trend toward more expressive input. For example, DesignPrompt [peng2024designprompt] allows designers to compose prompts using text, colour, and imagery, and DesignWeaver [tao2025designweaver] introduces palette-based refinement to support iterative design. Similarly, sketch-based interfaces such as Inkspire [lin2025inkspire] and SketchAI [davis2025sketchai] demonstrate how freehand input can guide image generation and support analogical inspiration.
@@ 2.1 Multimodal Human–AI Interaction for Creative Ideation
2.2 Talking as a Novel Input Modality for Sketching
Speech is also a natural modality for externalising thought, long recognised in systems such as SHRDLU [winograd1972understanding] and Put-That-There [bolt1980put] from the 1970-80s. It enables real-time, low-friction expression of ideas, making it especially valuable in early-stage design when thoughts are fluid and evolving [fry1979physics]. Sketching and speech together can serve as two intuitive, complementary interaction methods in design and creative domains, allowing users to express nuances that one modality alone might omit  [adler2007speech].
Recent works demonstrate how combining sketch and speech supports more natural and expressive communication: Rosenberg et al. [rosenberg2024drawtalking] introduced DrawTalking, which combines sketching with storytelling through speech to construct interactive animated worlds, while Giunchi et al. [giunchi2021mixing] showed that integrating sketch and speech for 3D model retrieval in virtual environments helped overcome the limitations of sketch-only input. Cheng et al. [cheng2025aiawareness] also explored speech and sketch as inputs for enhancing communication with AI through context awareness in an exploratory Wizard-of-Oz study.
@@ 2.2 Talking as a Novel Input Modality for Sketching
3 Formative Study: Understanding Early-Stage Design Workflow with GenAI
We first conducted a formative study, combining a design task with interviews, to understand how designers integrate GenAI into early-stage design and to identify opportunities for multimodal AI support to inform our subsequent system design in Section 4. Our institution’s ethics review board approved this study, and we obtained informed consent from all participants.
3.0.1 Participants
We recruited six participants (2 female, 4 male) with a basic background in design; at the very least, they had completed one design-related course (e.g., Urban Sketching or Design Thinking and Innovation). The participants included three students, two entry-level designers, and one experienced design practitioner, with design backgrounds spanning architecture, furniture, interior, robotics, and electronic product design. All participants had 2-5 years of part/full-time experience in design. They received approximately USD 7.8 as compensation for completing the 1-hour study.
3.0.2 Study Protocol
The study consists of three parts: pre-task interview (10 minutes), design task (30 minutes), and post-task interview (20 minutes). The pre-task interview captures participants’ design backgrounds, sketching habits, and prior experiences with GenAI tools. For the design task, we asked participants to design a household bread toaster and to express as many design ideas as possible under 30 minutes using a sketching application such as Goodnotes [^1] https://www.goodnotes.com/ or Procreate [^2] https://procreate.com/. They were permitted to utilize their personal devices–to ensure a familiar arrangement–or our Apple iPad Pro 13" and Apple Pencil set up with popular sketching apps (e.g., Goodnotes, Procreate) and GenAI tools (e.g., ChatGPT (GPT-4o) [^3] https://chatgpt.com/, Gemini [^4] https://gemini.google.com/app, Midjourney [^5] https://midjourney.com). The post-task interview then examined participants’ views on the AI’s role in supporting or limiting ideation, and captured their aspirations and expectations for an "ideal" AI assistant.
3.0.3 Data Analysis
We analysed the qualitative data by examining both participants’ interactions with sketching and AI tools during the design tasks and their reflections in post-study interviews. This allowed us to capture overall usage patterns, such as how participants sketched, engaged with GenAI tools, and iterated on ideas, while also identifying the challenges and gaps they experienced with current tools. These insights inform the design goals of our system.
3.1 Overall Usage Patterns and Challenges
@@
3.1.1 Pattern 1: Using GenAI for research and ideation.
Participants often used ChatGPT to explore design problems, define target users, or to inspire new directions. For instance, P1 asked about commonly found toaster issues and focused on cleaning functions, while P3 explored different user groups to inspire persona-driven designs. These uses supported functional exploration, but initial AI responses were often “too generic” (P3) and only became actionable after repeated prompt refinement.
3.1.2 Pattern 2: Using GenAI to render sketch-based ideas.
Several participants uploaded their sketches to ChatGPT Image or Gemini, or relied on text prompts, to visualise concepts. For instance, P2 uploaded a sideways-eject sketch, while P5 asked Gemini for a toaster with butter and egg compartments. However, the outputs often failed to match intent. P1 described the results as “kind of crazy,” P2 concluded, “I might have to redraw the whole thing so AI can understand,” and P5 added, “It takes too much time, I’d rather just draw.” Many participants ultimately returned to manual sketching for better clarity and control.
3.1.3 Pattern 3: Iterative loops across sketching, prompting, and referencing.
Participants often moved back and forth between tools. P3 only used text in ChatGPT to imagine playful forms (e.g., cat- or book-shaped toasters), while P1 first researched in ChatGPT, then sketched, went back to sketching after clarifications, and finally tried rendering images. P1 noted that “uploading sketches to ChatGPT was not so easy to operate on iPad.” This reflected how tool-switching slowed the process and introduced friction, especially for those who wanted to use image generation (P1, P2, P5, P6).
3.1.4 Challenges across patterns.
Across these usage patterns, participants encountered three main challenges: (1) AI responses that were too generic or required extensive refinement, (2) mismatched or low-quality image outputs that failed to convey intent, and (3) fragmented workflows due to frequent switching between different tools. These issues often interrupted the ideation flow and made participants rely more heavily on manual sketching for clarity and control.
3.2 Design Goals for TalkSketch
@@
3.2.1 Goal 1: Integrate AI with sketching tools.
@@
3.2.2 Goal 2: Reduce designers’ fatigue with long prompts.
@@
3.2.3 Goal 3: A more proactive and context-aware AI.
@@
4 TalkSketch: System Design and Development
@@
We first describe the technical implementation of TalkSketch (Figure 1), which integrates three core components: a digital sketching canvas for creating and managing drawings, to address the goal of integrating GenAI into sketching tools (Sec 3.2.1 Goal 1); a speech capture module that transcribes the user’s speech during sketching, aiming to support the goal of reducing fatigue from long prompts (Sec 3.2.2 Goal 2); and a multimodal AI chatbot that combines automatic insights with interactive text and image generation, aiming to make AI more proactive and context-aware (Sec 3.2.3 Goal 3). We then illustrate how these components work together through a system walkthrough using a concept toaster design example (Figure 2).
4.1 Technical Implementations
@@
TalkSketch integrates a unified multimodal AI chatbot that aims to support both proactive and user-initiated interaction. It comprises three complementary modules: Sketching, Talking, and Multimodal AI Chatbot.
4.1.1 Sketching Module
@@
The sketching canvas (Figure 1a) serves as the central workspace where participants draw early-stage design ideas using a stylus or touch input. Built using Fabric.js [^6] http://fabricjs.com/, the canvas supports drawing, erasing, selection, undo/redo, and canvas reset as illustrated in Figure 1. The top toolbar consolidates essential drawing controls along with the Generate with AI button. To facilitate interaction between users and AI, users can select any region of the sketching canvas and export it into the chatbot as part of a multimodal prompt. The Save to Gallery button allows users to save their canvas to the gallery once they are content with it; saved canvases can also be retrieved through the Gallery.
4.1.2 Talking Module
@@
TalkSketch augments the sketching experience with real-time voice capture to support think-aloud workflows. Audio recording initiates automatically when users are sketching (without the chatbot open), and the recording symbol will be turned on simultaneously to indicate this. Once the user opens the AI Chatbot, the recording stops and the captured audio is streamed for low-latency transcription via Google Cloud Speech-to-Text [^7] https://cloud.google.com/speech-to-text.
4.1.3 Multimodal AI Chatbot Module
@@
The multimodal AI chatbot (see Figures 1c and 1d) has two main components: (1) AI Insights, which automatically generate reflective feedback based on user sketches and verbal ideation, and (2) a Multimodal Chatbot Interface, where users can engage the AI through both text and sketch input. Both components share a common back-end powered by the Gemini series of multimodal models [^8] https://ai.google.dev/gemini-api/docs/models , known for their strong cross-modal reasoning capabilities and broad adoption across creative and analytical tasks [team2024gemini, yue2024mmmu]. Specifically, the text-based conversational features are supported by Gemini 2.0 Flash, while image generation is handled by Gemini 2.5 Flash Image. Both models accept multimodal inputs such as text and images, but they serve different purposes. Gemini 2.0 Flash is a fast multimodal-reasoning variant that produces text-only output, whereas Gemini 2.5 Flash Image is an image-generation variant of the Gemini 2.5 family, specialised for producing high-quality images from multimodal prompts. The Multimodal AI Chatbot Module maintains a unified conversation history that persists throughout the entire session and is consistently shared across both models, ensuring coherent context when switching between text and image generation.
@@
On the back-end, the AI Insights module is powered by Gemini 2.5 Flash and customised to simulate the role of a design thinking expert. Following the Double Diamond framework [tschimmel2012design], it guides users through the Discover and Define stages by identifying potential user needs, pain points, and framing design questions, before suggesting several exploratory directions to pursue. Two complementary prompt templates were developed to support different phases of the ideation process. When a user begins drawing on a new canvas, the system automatically triggers the Kickoff Prompt to initiate ideation; in subsequent interactions, as the user continues sketching or revising their transcript, the Refine Prompt is triggered to provide iterative feedback aligned with the evolving design concept. These prompts were iteratively tested to balance interpretability, creativity support, and brevity of response.
@@ 4.2 System Walkthrough
TalkSketch aims to support designers in early-stage ideation by combining freehand sketching, verbal thinking, and multimodal AI assistance. We illustrate its use through a walkthrough exemplar (Figure 2) where a fictional user Sky is tasked with designing a household toaster. Her interaction follows a natural flow: sketching while speaking aloud, triggering the chatbot interface, reviewing automatically generated AI insights, and exploring further via multimodal interaction.
4.2.1 Sketching while Talking.
Sky begins the design session by drawing a rough concept of a cube-shaped toaster with symbolic radial dials and a handle (see Figure 1a). As she sketches, she verbalises her thoughts, “I’m thinking of something bold and square, with a dial for heat control.” Her speech is recorded and transcribed in real time (see Figure 1b), forming a synchronised verbal-visual trace of her ideation process.
4.2.2 Entering the Chatbot and Viewing AI Insights.
When Sky clicks Generate with AI, the system enters the chatbot interface (see Figure 1d) and immediately displays an AI Insight (see Figure 1c). This insight is automatically generated from her sketch and think-aloud transcript, offering concise design reflections based on the Double Diamond design thinking framework. The system suggests emphasising geometric form or exploring tactile dials as directions worth pursuing. Sky reads the insight and finds it helpful in framing her design direction. She realises she had forgotten to mention a few details aloud–such as her intention for the toaster to have a retractable cord–and decides to edit the transcript directly. After adding the missing information, she regenerates the AI Insight to see how the suggestions change in response to the fuller context.
4.2.3 Exploring via Multimodal Interaction.
Sky then decides to follow up on the AI’s suggestions. She first types into the chatbot input box: “Could you give me some ideas for drawing a novel toaster?" In Text Generation Mode, the chatbot responds with directions such as transparent exteriors and interactive touchscreens. Sky incorporates several of these into her sketch. To visualise the design, she switches to Image Generation Mode, prompting: “Could you generate a realistic product based on my sketch?” The system returns a refined visual that preserves key features from her drawing.
@@
5 Potential Results
5.1 Enhancing Intent Expression and Interaction Naturalness